The mpfnmark Package^{∗}

W. Kowarschick

Technische Universiteat Muenchen Orleansstr. 34 81667 Muenchen

             kowa@informatik.tu-muenchen.de

1996/11/27

1 Introduction

The mpfnmark package provides the command \mpfootnotemark, which can be used in the same way as \footnotemark. The difference between these two macros is that within minipage environments the latter uses the standard footnote marker style (defined by \thefootnote), while the new command uses the minipage footnote marker style (defined by \thempfootnote).

2 A Sample Section Containing Only A Minipage

This is a sentence with a footnote inside a minipage.^{a} This is a sentence with a footnote outside that minipage.^{1} This is a sentence with the same footnote as the first sentence.^{a}

aThis is the footnote of the sentence with a footnote inside a minipage.

3 A Sample Section Containing No Minipage

This is a further sentence with a footnote outside the minipage.^{2} This is a sentence with the same footnote as the first sentence outside the minipage (marked by \footnotemark[2]).^{2} This is a further sentence with the same footnote as the first sentence outside the minipage (marked by \mpfootnotemark[2]).^{2}

∗This file has version number v1.0, last revised 1996/11/27, documentation dated 1996/11/27

1This is the footnote of the sentence with a footnote outside the minipage.

2This is the footnote of the further sentence with a footnote outside the minipage.

4 Source of the First Two Sections

\section{A Sample Section Containing Only A Minipage}
\hfill
\begin{minipage}{0.75\textwidth}
  This is a sentence with a footnote inside a minipage.\footnote{This is
    the footnote of the sentence with a footnote inside a minipage.} This is
  a sentence with a footnote outside that minipage.\footnotemark\ This is a
  sentence with the same footnote as the first sentence.\mpfootnotemark[1]\
\end{minipage}
\hfill
\footnotetext{
  This is the footnote of the sentence with a footnote outside the minipage.}
\section{A Sample Section Containing No Minipage}
This is a further sentence with a footnote outside the minipage.\footnote{
  This is the footnote of the further sentence with a footnote outside the
  minipage.} This is a sentence with the same footnote as the first sentence
outside the minipage (marked by |\footnotemark[2]|).\footnotemark[2]\
This is a further sentence with the same footnote as the first sentence
outside the minipage (marked by |\mpfootnotemark[2]|).\mpfootnotemark[2]\

5 The driver file

Nothing fancy.

1 ⟨∗driver⟩

2 \documentclass{ltxdoc}
3 \EnableCrossrefs

Some commonly used abbreviations for option names, filenames, counters and packages.

4 \newcommand\Lopt[1]{\textsf {#1}}
5 \newcommand\file[1]{\texttt {#1}}
6 \newcommand\Lcount[1]{\textsl {\small#1}}
7 \newcommand\pstyle[1]{\textsl {#1}}
8 \date{\docdate}

The body of the style. Include all the details.

9 \usepackage{mpfnmark}
10 \begin{document}
11 % \OnlyDescription %uncomment to suppress the source listing
12 \DocInput{mpfnmark.dtx}
13 \PrintIndex
14 \PrintChanges
15 \end{document}

16 ⟨/driver⟩

6 The package

17 ⟨∗package⟩

18 \NeedsTeXFormat{LaTeX2e}[1995/12/01]
19 \ProvidesPackage{mpfnmark}[1996/11/27 v1.0 W. Kowarschick LaTeX package]

20 \typeout{Package: `mpfnmark' v1.0\space<1996/11/27>}
21 \wlog{Documentation}

\mpfootnotemark This is the new macro. It is a modified version of \footnotemark found in latex.ltx distributed by the LaTeX3 project.

22 \def\mpfootnotemark{%
23 \@ifnextchar[\@xmpfootnotemark
24  {\stepcounter{footnote}%
25  \protected@xdef\@thefnmark{\thempfn}%
26  \@footnotemark}}
27 \def\@xmpfootnotemark[#1]{%
28 \begingroup
29  \c@footnote #1\relax
30  \unrestored@protected@xdef\@thefnmark{\thempfn}%
31 \endgroup
32 \@footnotemark}
33 % </package>

That’s us.

34 \endinput